OpenStudy (anonymous):

11. Use natural logarithms to solve the equation. Round to the nearest thousandth. 3e2x + 4 = 24 help!!

OpenStudy (whpalmer4):

\[3e^{2x}+4=24\]?

OpenStudy (anonymous):

subtract 4 divide by 3 take the log divide by 2

OpenStudy (anonymous):

when you take the log, use a calculator

OpenStudy (whpalmer4):

\[3e^{2x}+4=24\] \[3e^{2x}=20\]\[e^{2x}=\frac{20}{3}\]Take natural log of both sides to get rid of e^{2x}\[2x=\ln{\frac{20}{3}}\]\[x=\frac{1}{2}\ln{\frac{20}{3}}\]

OpenStudy (whpalmer4):

@satellite73 's steps are exactly right, just illustrating them so you can see why it works

OpenStudy (anonymous):

–0.960 0.499 1.117 0.949 but i dont get these answers?

OpenStudy (whpalmer4):

What do you get for the natural log of (20/3)?

OpenStudy (whpalmer4):

Are you perhaps getting 0.823...? That's log base 10

OpenStudy (anonymous):

i got 6.666637..

OpenStudy (whpalmer4):

Okay, 20/3 = 6.6666666... Now take the natural logarithm of that, and you should get 1.897... divide that by 2 and you get?

OpenStudy (anonymous):

.9485